OpenStudy (anonymous):

Solve 5^(x + 5) = 9^x

geerky42 (geerky42):

Take natural logarithm on both sides.

geerky42 (geerky42):

Can you handle it?

OpenStudy (anonymous):

ln 5^(x+5)= ln (9^x)

geerky42 (geerky42):

Yeah, do you know what to do with exponent?

OpenStudy (anonymous):

(x+5) ln(5)= ln (9^x)

geerky42 (geerky42):

Yes. Yes. Do the same thing on right side, then solve for x.

geerky42 (geerky42):

So you can find the x?

OpenStudy (anonymous):

here are the answer choices: x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69